Alice can vacuum in 40 minutes. It takes bob 45 minutes to do the same job. Find the time it takes Alice and Bob to do the job together.
Answer by venugopalramana(3286)

: If it takes me 4 hours to paint a huse and it take you 6 hours, how long willit take both of us working together?1 solutions Answer 9497 by venugopalramana(345) on 2005-11-15 10:54:53  SEE THE FOLLOWING AND TRY YOUR PROBLEM.IF THERE IS STILL DIFFICULTY PLEAS COME BACK Answer 9266 by venugopalramana(334) on 2005-11-11 01:54:43  SEE THIS EXAMPLE AND DO IN THE SAME MANNER Joe digs a ditch in 3.2 hours.Joe working with George can did the same length ditch in 2.7 hours.How long would it take George to dig the ditch by himself. 1 solutions --------------------------------------------------------------------------------Answer 9228 by venugopalramana(299) on 2005-11-10 06:00:00  Joe digs a ditch in 3.2 hours.Joe working with George can did the same length ditch in 2.7 hours.How long would it take George to dig the ditch by himself. Joe digs a ditch in 3.2 hours...so in 1 hour he digs 1/3.2 ditch=10/32=5/16let george alone take x hrs to dig the ditch...so in 1 hour he digs 1/x ditchso in 1 hour joe and george together dig 5/16+1/x ditch =(5x+16)/16x so they take 16x/(5x+16) hrs to dig the ditch togetherbut this given as 2.7 hoursso 16x/(5x+16)=2.7cross multiplying we get 16x=2.7(5x+16)=13.5x+43.216x-13.5x=43.22.5x=43.2x=43.2/2.5=432/25=17.28 hoursso george alone takes 17.28 hrs to dig the ditch